What is the coefficient of x^3 y^13 in the expansion of (−1x+2y)^16?
1
Expert's answer
2020-11-09T15:30:38-0500

By Binomial Theorem


"(a+b)^n=\dbinom{n}{0}a^nb^0+\dbinom{n}{1}a^{n-1}b^1+...+"

"+\dbinom{n}{k}a^kb^{n-k}+...+\dbinom{n}{n}a^0b^n"

We have "a=-x, b=2y, n=16."

We find the coefficient of "x^3 y^{13}"


"k=3, \dbinom{n}{k}a^kb^{n-k}=\dbinom{16}{3}(-x)^3(2y)^{16-3}="

"=-\dfrac{16!}{3!(16-3)!}(8192)x^3y^{13}=-560(8192)x^3y^{13}="

"=-4587520x^3y^{13}"

The coefficient of "x^3y^{3}" in the expansion of "(-1x+2y)^{16}" is "-4587520."